OpenStudy (daniel.ohearn1):

Help Please. What is the integral of (x-9)/((x+5)(x-2))dx?

OpenStudy (freckles):

have you tried partial fractions?

OpenStudy (daniel.ohearn1):

Yes

OpenStudy (freckles):

\[\frac{x-9}{(x+5)(x-2)}=\frac{A}{x+5}+\frac{B}{x-2}\]

OpenStudy (freckles):

your answer will contain natural log

OpenStudy (freckles):

after the integration

OpenStudy (freckles):

what did you get for A and B?

OpenStudy (daniel.ohearn1):

I got confusion

OpenStudy (freckles):

can you show me your attempt and maybe we can fix it together

OpenStudy (daniel.ohearn1):

A(x-2) + B(x+5)= x-9

OpenStudy (freckles):

that's good

OpenStudy (freckles):

now your can plug in values of x to find A and B or you can regroup things and compare both sides to find A and B which sounds more fun to you?

OpenStudy (daniel.ohearn1):

I have to go for a bit... I've attempted it but I'm going to college now my ride is arriving...

OpenStudy (daniel.ohearn1):

afk tyyl

OpenStudy (freckles):

k